Calculate the acceleration of the bob of the simple pendulum of time period 2s when the displacement from the equilibrium position is 42 mm 


1
Expert's answer
2021-02-24T12:50:10-0500
"a=A\omega^2=A(\dfrac{2\pi}{T})^2,""a=42\cdot10^{-3}\ m\cdot(\dfrac{2\pi}{2\ s})^2=0.414\ \dfrac{m}{s^2}."